OpenStudy (anonymous):

solve the equation: √x+6+√2-x=4

myininaya (myininaya):

\[\sqrt{x+6}+\sqrt{2-x}=4?\]

OpenStudy (anonymous):

yeah, that's my question

myininaya (myininaya):

\[\sqrt{x+6}=4-\sqrt{2-x}\]

OpenStudy (anonymous):

i don't know how to type it just like that in the posts..

myininaya (myininaya):

first thing i did was subracted sqrt(2-x) on both sides

myininaya (myininaya):

now i'm going to square both sides

myininaya (myininaya):

remember \[(a+b)^2=a^2+2ab+b^2\] \[(\sqrt{x+6})^2=(4-\sqrt{2-x})^2\] \[x+6=15-2(4)(\sqrt{2-x})+(\sqrt{2-x})^2\]

myininaya (myininaya):

oops 4^2 is 16

myininaya (myininaya):

type-o

myininaya (myininaya):

\[x+6=16-2(4)\sqrt{2-x}+(\sqrt{2-x})^2\]

myininaya (myininaya):

\[x+6=16-8\sqrt{2-x}+(2-x)\]

myininaya (myininaya):

now we simplify and put sqrt{2-x} by itself on one side \[x+6=18-x-8\sqrt{2-x}\] \[x+x+6-18=-8\sqrt{2-x}\]

myininaya (myininaya):

\[2x-12=-8\sqrt{2-x}\] we can leave the -8 there in front of the sqrt{2-x} now square both sides \[(2x-12)^2=(-8\sqrt{2-x})^2\]

OpenStudy (anonymous):

lol

myininaya (myininaya):

\[4x^2-2(2)(12)x+144=64(2-x)\] \[4x^2-48x+144=128-64x\] see we don't have any square root thingy ma jiggers to worry about now

myininaya (myininaya):

now we put everything on one side and see if we can factor or we can use the quadratic formula

myininaya (myininaya):

\[4x^2-48x+64x+144-128=0\]

myininaya (myininaya):

\[4x^2+16x+ 16=0\]

myininaya (myininaya):

now each term has the factor 4 in common we can factor 4 out and then divide both sides by 4 \[x^2+4x+4=0\]

myininaya (myininaya):

this can easily be factored (x+2)^2=0

myininaya (myininaya):

x+2=0 x=-2

OpenStudy (anonymous):

that seemed like a lot of work and you just whipped it out like nothing. THANK you, first person where they actually care about helping and explaining

myininaya (myininaya):

i don't see the point in just giving an answer without an explanation

OpenStudy (anonymous):

i very much appreciate that!! you have NO idea

myininaya (myininaya):

np